OpenStudy (anonymous):

cos(-247.5 DEGREES) IN EXACT VALUE. solve using half angle formula.

OpenStudy (inkyvoyd):

Hm.

OpenStudy (inkyvoyd):

add 360 degrees.

OpenStudy (inkyvoyd):

tell me what you get.

OpenStudy (anonymous):

112.5

OpenStudy (inkyvoyd):

Now double that value!

OpenStudy (anonymous):

ohh so cos 225 - sqroot 2 over 2?

OpenStudy (inkyvoyd):

You know the half angle formula?

OpenStudy (inkyvoyd):

I forgot it, but I can derive it on spot :D

OpenStudy (anonymous):

yes cos u/2 = sqroot 1 + cos u/2

OpenStudy (inkyvoyd):

Yep.

OpenStudy (anonymous):

http://mathworld.wolfram.com/Half-AngleFormulas.html

OpenStudy (inkyvoyd):

So, you should be fine. If you want to check your work, go to wolfram alpha, and just type in the cos(blah)

OpenStudy (anonymous):

I have tried that on both wolfram and my calculator I get an exact amount

OpenStudy (inkyvoyd):

You get an exact amount, or you want one but they dont' give one?

OpenStudy (anonymous):

i got the problem wrong... i got sqroot 2+sqroot2 all over 2

OpenStudy (anonymous):

but it's wrong

OpenStudy (inkyvoyd):

Ok, lemme do some math

OpenStudy (inkyvoyd):

That's probaly a quadrant eror.

OpenStudy (inkyvoyd):

And, @lgbasallote

OpenStudy (lgbasallote):

someone called?

OpenStudy (inkyvoyd):

Helpy out?

OpenStudy (lgbasallote):

hmmm this seems difficult -__-- that or i just forgot my trig

OpenStudy (inkyvoyd):

Add 360 derees, then use 225 as the angle to find half of.

OpenStudy (inkyvoyd):

-247.5 bro.

OpenStudy (anonymous):

\[ t =247.5 \\ 2t= 495= 360 + 90 +45 \\ \cos(2t) =\ cos(90 + 45)=-\cos(45) = -\frac 1{\sqrt 2}\\ \cos^2(t)=\frac{ 1+ \cos(2t)}2=\frac {1-\frac 1{\sqrt 2}} {2}=\\ \frac{1}{2}-\frac{1}{2 \sqrt{2}}\\ \cos(t) = \pm \sqrt{\frac{1}{2}-\frac{1}{2 \sqrt{2}}} \]

OpenStudy (anonymous):

\[ \cos(t) = \cos(247.5)= -\sqrt{\frac{1}{2}-\frac{1}{2 \sqrt{2}}}= -0.382683 \]